OpenStudy (gabylovesyou):

Pablo rolled a standard die 30 times. He got a 1 eight times. How does this result compare to the expected results? Explain. Please explain (confused) :S

OpenStudy (gabylovesyou):

@dpaInc Need ur help!! lol

jimthompson5910 (jim_thompson5910):

A standard die has 6 sides. You expect each side to have an equal chance of landing face up. So the probability of getting any face is 1/6 So the probability of getting a 1 showing face up is 1/6 Out of 30 times, you'll land a "1" 30(1/6) = 30/6 = 5 times So Pablo rolled a "1" three more times than expected.

OpenStudy (gabylovesyou):

why 3 more times??

jimthompson5910 (jim_thompson5910):

because we expect to roll a "1" five times we expect to roll a "2" five times ... etc for all 6 numbers So each number should show up 5 times. There are 6 numbers, which totals to 5*6 = 30 rolls.

OpenStudy (gabylovesyou):

ohhhhh ok!

jimthompson5910 (jim_thompson5910):

But he rolled a "1" 8 times, which is 8-5 = 3 more than expected It's not like it's impossible, it's just that the average number of rolls should be 5.

OpenStudy (gabylovesyou):

ok! perfectoo lol

jimthompson5910 (jim_thompson5910):

lol awesome